OpenStudy (anonymous):

As first public inquiry showed, there were 627,000 women motorcyclists in the United States, up from 447,000 just eight years before it. Find the percent of increase in the number of women motorcyclists during these eight years. The Percent of increase is___% (Round too the nearest Tenth as needed)

OpenStudy (anonymous):

do i round off 447000 too 500000

OpenStudy (calculusfunctions):

change in % = (new amount - original amount)/(original amount) If this change is positive, then an increase in percent occurs. If the change is negative, then a & decrease occurs. Thus in your question, there will be a percent increase because the new figure is larger than the original figure. Thus \[% increase =\frac{ 627 000-447 000 }{447 000 }\] \[=\frac{ 180 000 }{ 447 000 }\] \[=0.403\] ∴ there is about a 40% increase. Does that help?

OpenStudy (anonymous):

yes. would i have too round 0.403 too the nearest tenth

OpenStudy (calculusfunctions):

No, rounding to 40% as I did is fine.

OpenStudy (calculusfunctions):

Don't I get a medal?

OpenStudy (anonymous):

how do i give u a medal

OpenStudy (calculusfunctions):

You just did, I think. lol. Thanks!

OpenStudy (anonymous):

40% is the answer, Right

OpenStudy (calculusfunctions):

Do you understand now, how to arrive at that answer?

OpenStudy (anonymous):

iom still confuse.lol

OpenStudy (calculusfunctions):

Did you read my explanation at the beginning of the solution I provided.

OpenStudy (anonymous):

yes.

OpenStudy (calculusfunctions):

Tell me where you're confused, and perhaps I can help clear the doubts.

OpenStudy (anonymous):

Im jsut confuse with the percent of increase.

OpenStudy (calculusfunctions):

If the new amount is larger than the original, then there will be a percent increase. If the new amount is smaller than the original amount, then there will be a percent decrease. To work it out, I gave you the formula at the beginning of the solution I did.

OpenStudy (anonymous):

=180000 447000 =0.403 Do you have too change 0.403 into a percent

OpenStudy (calculusfunctions):

Sorry, I just logged back in to the site. Yes, you should express your final answer in terms of percent.